OpenStudy (anonymous):

Find the quadratic equation whose solutions are: 2 + sqr(5) and 2 - sqr(5) I'm not just looking for an answer but rather to learn.

OpenStudy (anonymous):

\[2+ \sqrt{5} and 2 - \sqrt{5}\]

OpenStudy (anonymous):

equation bcam x^2 - (a+b)x+ a.b =0 take a=2+sqr 5 , b=2-sqr -5

OpenStudy (anonymous):

\[x ^{2} - (a+b)x + ab = 0?\]

OpenStudy (anonymous):

this is from an example in my book, but I don't understand the example:

OpenStudy (anonymous):

tis what any quarratic equation can be form...

OpenStudy (anonymous):

if their two roots are given .got it ...now

OpenStudy (anonymous):